# 3D Diffuser Actor: Multi-task 3D Robot Manipulation with Iterative Error Feedback

### v2

@@ IV-A Evaluation on RLBench
Avg. open slide sweep to meat off turn put in close drag stack Success $\uparrow$ drawer block dustpan grill tap drawer jar stick blocks C2F-ARM-BC [33] 20.1 20 16 0 20 68 4 24 24 0 PerAct [69] 49.4 88.0±${}_{\pm 5.7}$ 74.0±${}_{\pm 13.0}$ 52.0±${}_{\pm 0.0}$ 70.4±${}_{\pm 2.0}$ 88.0±${}_{\pm 4.4}$ 51.2±${}_{\pm 4.7}$ 55.2±${}_{\pm 4.7}$ 89.6±${}_{\pm 4.1}$ 26.4±${}_{\pm 3.2}$ HiveFormer [23] 45 52 64 28 100 80 68 52 76 8 PolarNet [8] 46 84 56 52 100 80 32 36 92 4 RVT [22] 62.9 71.2±${}_{\pm 6.9}$ 81.6±${}_{\pm 5.4}$ 72.0±${}_{\pm 0.0}$ 88.0±${}_{\pm 2.5}$ 93.6±${}_{\pm 4.1}$ 88.0±${}_{\pm 5.7}$ ${}_{\pm 2.5}$ 99.2±${}_{\pm 1.6}$ 28.8±${}_{\pm 3.9}$ Act3D [19] 63.2 78.4±${}_{\pm 11.2}$ 96.0±${}_{\pm 2.5}$ 86.4±${}_{\pm 6.5}$ 95.2±${}_{\pm 1.6}$ 94.4±${}_{\pm 2.0}$ 91.2±${}_{\pm 6.9}$ 96.8±${}_{\pm 3.0}$ 80.8±${}_{\pm 6.4}$ 4.0±${}_{\pm 3.6}$ 3D Diffuser Actor (ours) 81.3 89.6±${}_{\pm 4.1}$ 97.6±${}_{\pm 3.2}$ 84.0±${}_{\pm 4.4}$ 96.8±${}_{\pm 1.6}$ 99.2±${}_{\pm 1.6}$ 96.0±${}_{\pm 3.6}$ 96.0±${}_{\pm 2.5}$ 100.0±${}_{\pm 0.0}$ 68.3±${}_{\pm 3.3}$ screw put in place put in sort push insert stack place bulb safe wine cupboard shape buttons peg cups cups C2F-ARM-BC [33] 8 12 8 0 8 72 4 0 0 PerAct [69] 17.6±${}_{\pm 2.0}$ 86.0±${}_{\pm 3.6}$ 44.8±${}_{\pm 7.8}$ 28.0±${}_{\pm 4.4}$ 16.8±${}_{\pm 4.7}$ 92.8±${}_{\pm 3.0}$ 5.6±${}_{\pm 4.1}$ 2.4±${}_{\pm 2.2}$ 2.4±${}_{\pm 3.2}$ HiveFormer [23] 8 76 80 32 8 84 0 0 0 PolarNet [8] 44 84 40 12 12 96 4 8 0 RVT [22] 48.0±${}_{\pm 5.7}$ 91.2±${}_{\pm 3.0}$ 91.0±${}_{\pm 5.2}$ 49.6±${}_{\pm 3.2}$ 36.0±${}_{\pm 2.5}$ 100.0±${}_{\pm 0.0}$ 11.2±${}_{\pm 3.0}$ 26.4±${}_{\pm 8.2}$ 4.0±${}_{\pm 2.5}$ Act3D [19] 32.8±${}_{\pm 6.9}$ 95.2±${}_{\pm 4.0}$ 59.2±${}_{\pm 9.3}$ 67.2±${}_{\pm 3.0}$ 29.6±${}_{\pm 3.2}$ 93.6±${}_{\pm 2.0}$ 24.0±${}_{\pm 8.4}$ 9.6±${}_{\pm 6.0}$ 3.2±${}_{\pm 3.0}$ 3D Diffuser Actor (ours) 82.4±${}_{\pm 2.0}$ 97.6±${}_{\pm 2.0}$ 93.6±${}_{\pm 4.8}$ 85.6±${}_{\pm 4.1}$ ${}_{\pm 4.4}$ 98.4±${}_{\pm 2.0}$ 65.6±${}_{\pm 4.1}$ 47.2±${}_{\pm 8.5}$ 24.0±${}_{\pm 7.6}$
@@
TABLE I: Multi-Task performance on RLBench. We report success rates on 18 RLBench tasks of the same test split as PerAct [69]. Four camera views are used for all experiments. We evaluate the final checkpoint over 5 random seeds with the same 3D object models and success conditions of RLBench [32] as PerAct. We show the mean and standard deviation of success rates average across all random seeds. Our 3D Diffuser Actor outperforms all prior arts among most tasks by a large margin. Variances are included when available.
Results on the PerAct setup We show quantitative results in Table I. Our 3D Diffuser Actor outperforms all baselines on most tasks by a large margin. It achieves an average $81.3\%$ success rate among all 18 tasks, an absolute improvement of $+18.1\%$ over Act3D, the previous state-of-the-art. In particular, 3D Diffuser Actor achieves big leaps on tasks with multiple modes, such as stack blocks, stack cups and place cups, which most baselines fail to complete. We obtain substantial improvements of $+39.5\%$, $+18.4\%$, $+41.6\%$ and $+20.8\%$ on stack blocks, put in cupboard, insert peg and stack cups respectively.
@@
Avg. close open sweep to meat off turn slide put in drag push stack Success. jar drawer dustpan grill tap block drawer stick buttons blocks GNFactor [83] 31.7 25.3 76.0 28.0 57.3 50.7 20.0 0.0 37.3 18.7 4.0 Act3D [19] 65.3 52.0 84.0 80.0 66.7 64.0 100.0 54.7 86.7 64.0 0.0 3D Diffuser Actor 78.4 82.7 89.3 94.7 88.0 80.0 92.0 77.3 98.7 69.3 12.0
TABLE II: Multi-Task performance on RLBench with single camera view. Following GNFactor [83], we report sucess rates on 10 RLBench tasks using only the front camera view. All models are trained with 20 demonstrations per task. We evaluate the final checkpoints across 3 seeds with 25 episodes for each task in the test set. Our 3D Diffuser Actor outperforms prior state-of-the-art baselines–GNFactor and Act3D–among most tasks by a large margin.
Results on the GNFactor setup We show quantitative results in Table II. We train Act3D on this setup using its publicly available code. 3D Diffuser Actor outperforms both GNFactor and Act3D by a significant margin, achieving absolute performance gains of $+46.4\%$ and $+13.1\%$ respectively. Notably, 3D Diffuser Actor and Act3D utilize similar 3D scene representations—sparse 3D feature tokens—while GNFactor featurizes a scene with 3D voxels and learns to reconstruct them. Even with a single camera view, both Act3D and 3D Diffuser Actor outperform GNFactor significantly. This suggests that the choice of 3D scene representation is a more crucial factor than the completion of 3D scenes in developing efficient 3D manipulation policies.
Avg. Success. 2D Diffuser Actor 47.0 3D Diffuser Actor w/o Rel. Attn. 71.3 3D Diffuser Actor (ours) 81.3
@@
TABLE III: Ablation study. We follow the same simulation setup as [69] and report average success rate on 18 RLBench tasks. 3D Diffuser Actor significantly outperforms its counterparts that do not use 3D scene representations or translation-equivariant 3D relative attention.
@@
Ablations We ablate the use of 3D scene representations and the use of relative attentions in the PerAct experimental setup in Table III. We draw the following conclusions: 1. 3D Diffuser Actor largely outperforms its 2D counterpart, 2D Diffuser Actor, underlining the importance of 3D scene representations. 2. 3D Diffuser Actor with absolute 3D attentions (3D Diffuser Actor w/o Rel. Attn.) is much worse than 3D Diffuser Actor with relative 3D attentions. This shows that translation equivariance through relative attentions is very important for generalization. Notably, this baseline already outperforms all prior arts in Table I, proving the effectiveness of marrying 3D representations and diffusion policies.
Train Task completed in a row episodes 1 2 3 4 5 Avg. Len MCIL [44] All 30.4 1.3 0.2 0.0 0.0 0.31 HULC [48] All 41.8 16.5 5.7 1.9 1.1 0.67 RT-1 [4] Lang 53.3 22.2 9.4 3.8 1.3 0.90 RoboFlamingo [38] Lang 82.4 61.9 46.6 33.1 23.5 2.48 SuSIE [2] All 87.0 69.0 49.0 38.0 26.0 2.69 GR-1 [80] Lang 85.4 71.2 59.6 49.7 40.1 3.06 3D Diffuser Actor (ours, 60 keyposes) Lang 92.2 78.7 63.9 51.2 41.2 3.27 3D Diffuser Actor (ours, 360 keyposes) Lang 95.7 88.1 78.6 65.7 55.3 3.83
TABLE IV: Zero-shot long-horizon performance on CALVIN. We report success rates and average length of completed task sequences. All models are trained on play trajectories on environments A, B, C, and tested on 1000 unique instruction chains on environment D. MCIL, HULC and SuSIE are trained with all play trajectories in the training scenes, whereas, RT-1, RoboFlamingo, GR-1 and our 3D Diffuser Actor are trained with the subset of play data annotated with language descriptions. 3D Diffuser Actor achieves state-of-the-art performance. For reference, we include our model’s performance when allowed for more predictions, in which case the performance significantly increases.
IV-B Evaluation on CALVIN
@@
The CALVIN benchmark is build on top of the PyBullet [13] simulator and involves a Franka Panda Robot arm that manipulates the scene. CALVIN consists of 34 tasksThe definition of “task” varies between different benchmarks: CALVIN considers the manipulation of variations (e.g., same objects with different colors) as different tasks, whereas, RLBench considers those the same task. and 4 different environments (A, B, C and D). All environments are equipped with a desk, a sliding door, a drawer, a button that turns on/off an LED, a switch that controls a lightbulb and three different colored blocks (red, blue and pink). These environments differ from each other in the texture of the desk and positions of the objects. CALVIN provides 24 hours of tele-operated unstructured play data, 35% of which are annotated with language descriptions.
We train 3D Diffuser Actor on the subset of play data annotated with language descriptions on the environments A, B and C and evaluate it on 1000 unique instruction chains on environment D, following prior works  [38, 80]. Each instruction chain includes five language instructions that need to be executed sequentially. We devise an algorithm (Appendix, Section A-A) to extract keyposes on CALVIN, since prior works do not use keyposes on this benchmark. We train our 3D Diffuser Actor to predict both the end-effector pose for each keypose and the corresponding trajectory to reach the predicted pose, instead of using a motion planner. Prior works [48, 2, 38] predict a maximum of 360 actions to complete each instructional task, while, on average, it takes only 60 actions to complete each task using ground-truth trajectories. Our model predicts both keyposes and corresponding trajectories, and, on average, it takes 10 keyposes of each demonstration to complete an instructional task. We thus allow our model to predict a maximum of 60 keyposes for each task. For reference, we also allow our model to predict 360 times and show the influence of this hyperparameter in the Appendix (Section A-B).
Baselines. We consider the following baselines:
  1. 1.
    MCIL [44], a multi-modal goal-conditioned 2D policy that maps three types of goals–goal images, language instructions and task labels–to a shared latent feature space, and conditions on such latent goals to predict actions.
  2. 2.
    HULC [48], a 2D policy that uses a variational autoencoder to sample a latent plan based on the current observation and task description, then conditions on this latent to predict actions.
  3. 3.
    RT-1 [4], a 2D transformer-based policy that encodes the image and language into a sequence of tokens and employs a Transformer-based architecture that contextualizes these tokens and predicts the arm movement or terminates the episode.
  4. 4.
    RoboFlamingo [38], a 2D policy that adapts existing vision-language models, which are pre-trained for solving vision and language tasks, to robot control. It uses frozen vision and language foundational models and learns a cross-attention between language and visual features, as well as a recurrent policy that predicts the low-level actions conditioned on the language latents.
  5. 5.
    SuSIE [2], a 2D policy that deploys an large-scale pre-trained image generative model [6] to synthesize visual subgoals based on the current observation and language instruction. Actions are then predicted by a low-level goal-conditioned 2D diffusion policy that models inverse dynamics between the current observation and the predicted subgoal image.
  6. 6.
    GR-1 [80], a 2D policy that first pre-trains an autoregressive Transformer on next frame prediction, using a large-scale video corpus without action annotations. Each video frame is encoded into an 1d vector by average-pooling its visual features. Then, the same architecture is fine-tuned in-domain to predict both actions and future observations.
We report results for HULC, RoboFlamingo, SuSIE and GR-1 from the respective papers. Results from MCIL are borrowed from [48]. Results from RT-1 are copied from [80].
Evaluation metrics Following previous works [48, 80], we report the success rate and the average number of completed sequential tasks.
Results We present our results in Table IV. 3D Diffuser Actor achieves competitive results compared to the state-of-the-art models. In comparison to GR-1, our model achieves absolute performance gains of $+6.8\%$, $+7.5\%$, $+4.3\%,+1.5\%$ and $+1.1\%$ for completing the five tasks in the instruction chain. On average, 3D Diffuser Actor complete $3.27$ tasks in a row, setting a new state of the art. Note that GR-1’s contribution is large-scale pre-training on videos, which is a scheme orthogonal to our contribution. Lastly, allowing our model to predict more times increases the performance significantly. For this performance, we found stronger conditioning on the language instructions to be important, as we explain in our appendix (Figure 6b).
@@
TABLE V: Multi-Task performance on real-world tasks.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption "insert the peg "put grapes "put a mouse into a hole" in the bowl" on the pad"
Figure 4: Visualized results of our real-world manipulation.
IV-C Evaluation in the real world
We validate 3D Diffuser Actor in learning manipulation tasks from real-world demonstrations. We use a Franka Emika robot and capture visual observations with a Azure Kinect RGB-D sensor at a front view. Images are originally captured at $1280\times 720$ resolution and downsampled to a resolution of $256\times 256$. Camera extrinsics are calibrated w.r.t the robot base. We use 12 tasks: 1) close a box, 2) put ducks in bowls, 3) insert a peg vertically into the hole, 4) insert a peg horizontally into the torus, 5) put a computer mouse on the pad, 6) open the pen, 7) press the stapler, 8) put grapes in the bowl, 9) sort the rectangle, 10) stack blocks with the same shape, 11) stack cups and 12) put block in a triangle on the plate. Details on the task definitions and success criteria are included in the Appendix (Section A-C).

### v3

@@ 4.1 Evaluation on RLBench
Avg. open slide sweep to meat off turn put in close drag stack Success $\uparrow$ drawer block dustpan grill tap drawer jar stick blocks C2F-ARM-BC [james2022coarse] 20.1 20 16 0 20 68 4 24 24 0 PerAct [shridhar2023perceiver] 49.4 88.0±5.7 74.0±13.0 52.0±0.0 70.4±2.0 88.0±4.4 51.2±4.7 55.2±4.7 89.6±4.1 26.4±3.2 HiveFormer [guhur2023instruction] 45 52 64 28 100 80 68 52 76 8 PolarNet [Chen2023PolarNet3P] 46 84 56 52 100 80 32 36 92 4 RVT [goyal2023rvt] 62.9 71.2±6.9 81.6±5.4 72.0±0.0 88.0±2.5 93.6±4.1 88.0±5.7 52.0±2.5 99.2±1.6 28.8±3.9 Act3D [gervet2023act3d] 63.2 78.4±11.2 96.0±2.5 86.4±6.5 95.2±1.6 94.4±2.0 91.2±6.9 96.8±3.0 80.8±6.4 4.0±3.6 3D Diffuser Actor (ours) 81.3 89.6±4.1 97.6±3.2 84.0±4.4 96.8±1.6 99.2±1.6 96.0±3.6 96.0±2.5 100.0±0.0 68.3±3.3 screw put in place put in sort push insert stack place bulb safe wine cupboard shape buttons peg cups cups C2F-ARM-BC [james2022coarse] 8 12 8 0 8 72 4 0 0 PerAct [shridhar2023perceiver] 17.6±2.0 86.0±3.6 44.8±7.8 28.0±4.4 16.8±4.7 92.8±3.0 5.6±4.1 2.4±2.2 2.4±3.2 HiveFormer [guhur2023instruction] 8 76 80 32 8 84 0 0 0 PolarNet [Chen2023PolarNet3P] 44 84 40 12 12 96 4 8 0 RVT [goyal2023rvt] 48.0±5.7 91.2±3.0 91.0±5.2 49.6±3.2 36.0±2.5 100.0±0.0 11.2±3.0 26.4±8.2 4.0±2.5 Act3D [gervet2023act3d] 32.8±6.9 95.2±4.0 59.2±9.3 67.2±3.0 29.6±3.2 93.6±2.0 24.0±8.4 9.6±6.0 3.2±3.0 3D Diffuser Actor (ours) 82.4±2.0 97.6±2.0 93.6±4.8 85.6±4.1 44.0±4.4 98.4±2.0 65.6±4.1 47.2±8.5 24.0±7.6
@@
Table 1: Evaluation on RLBench on the multi-view setup. We show the mean and standard deviation of success rates average across all random seeds. 3D Diffuser Actor outperforms all prior arts on most tasks by a large margin. Variances are included when available.
@@
Avg. close open sweep to meat off turn slide put in drag push stack Success $\uparrow$ jar drawer dustpan grill tap block drawer stick buttons blocks GNFactor [ze2023gnfactor] 31.7 25.3 76.0 28.0 57.3 50.7 20.0 0.0 37.3 18.7 4.0 Act3D [gervet2023act3d] 65.3 52.0±5.7 84.0±8.6 80.0±9.8 66.7±1.9 64.0±5.7 100.0±0.0 54.7±3.8 86.7±1.9 64.0±1.9 0.0±0.0 3D Diffuser Actor 78.4 82.7±1.9 89.3±7.5 94.7±1.9 88.0±5.7 80.0±8.6 92.0±0.0 77.3±3.8 98.7±1.9 69.3±5.0 12.0±3.7
Table 2: Evaluation on RLBench on the single-view setup. 3D Diffuser Actor outperforms prior state-of-the-art baselines, GNFactor and Act3D, on most tasks by a large margin.
We show quantitative results for the multi-view setup in Table 1 and for single-view setup in Table 2. On multi-view, 3D Diffuser Actor achieves an average $81.3\%$ success rate among all 18 tasks, an absolute improvement of $+18.1\%$ over Act3D, the previous state-of-the-art. In particular, 3D Diffuser Actor achieves big leaps on long-horizon high-precision tasks with multiple modes, such as stack blocks, stack cups and place cups, which most baselines fail to complete. All baselines use classification or regression losses, our model is the only one to use diffusion for action prediction.
On single-view, 3D Diffuser Actor outperforms GNFactor by $+46.7\%$ and Act3D by $+13.1\%$. Surprisingly, Act3D also outperforms GNFactor by a big margin. This suggests that the choice of 3D scene representation is more crucial than 3D feature completion. Since Act3D has a similar 3D scene tokenization as 3D Diffuser Actor, this shows the importance of diffusion over alternatives to handle multimodality in prediction, specifically over sampling based 3D action maps and rotation regression.
Ablations We consider the following ablative versions of our model: 1. 2D Diffuser Actor, our implementation of 2D Diffusion Policy of [chi2023diffusion]. We remove the 3D scene encoding from 3D Diffuser Actor and instead use per-image 2D representations by average-pooling features within each view. We add learnable embeddings to distinguish between different views and fuse them with action estimates, as done in [chi2023diffusion]. 2. 3D Diffuser Actor w/o RelAttn., an ablative version that uses absolute (non-relative) attentions.
Avg. Success. 2D Diffuser Actor 47.0 3D Diffuser Actor w/o RelAttn. 71.3 3D Diffuser Actor (ours) 81.3
@@
Table 3: Ablation study. Our model significantly outperforms its counterparts that do not use 3D scene representations or translation-equivariant 3D relative attention.
@@
We show ablative results in Table 3. 3D Diffuser Actor largely outperforms its 2D counterpart, 2D Diffuser Actor. This shows the importance of 3D scene representations in performance. 3D Diffuser Actor with absolute 3D attentions (3D Diffuser Actor w/o RelAttn.) is worse than 3D Diffuser Actor with relative 3D attentions. This shows that translation equivariance through relative attentions is important for generalization. Despite that, this ablative version already outperforms all prior arts in Table 1, proving the effectiveness of marrying 3D representations and diffusion policies.
4.2 Evaluation on CALVIN
@@
The CALVIN benchmark is built on top of the PyBullet [coumans2016pybullet] simulator and involves a Franka Panda Robot arm that manipulates the scene. CALVIN consists of 34 tasks and 4 different environments (A, B, C and D). All environments are equipped with a desk, a sliding door, a drawer, a button that turns on/off an LED, a switch that controls a lightbulb and three different colored blocks (red, blue and pink). These environments differ from each other in the texture of the desk and positions of the objects. CALVIN provides 24 hours of tele-operated unstructured play data, 35% of which are annotated with language descriptions. Each instruction chain includes five language instructions that need to be executed sequentially. We evaluate on the so called zero-shot generalization setup, where models are trained in environments A, B and C and tested in D. We report the success rate and the average number of completed sequential tasks, following previous works [mees2022hulc, wu2023unleashing]. No motion planner is available in CALVIN so all models need to predict robot pose trajectories.
Baselines All methods tested so far in CALVIN are 2D policies, that do not use depth or camera extrinsics. We compare against the hierarchical 2D policies of MCIL [lynch2020language], HULC [mees2022hulc] and SuSIE [black2023zero] which predict latent features or images of subgoals given a language instruction, which they feed into lower-level subgoal-conditioned policies. They can train the low-level policy on all data available in CALVIN as opposed to the language annotated subset only. We compare against large scale 2D transformer policies of RT-1 [brohan2022rt], RoboFlamingo [li2023vision] and GR-1 [wu2023unleashing] which pretrain on large amounts of interaction or observational (video alone) data. We report results for HULC, RoboFlamingo, SuSIE and GR-1 from the respective papers. Results from MCIL are borrowed from [mees2022hulc]. Results from RT-1 are copied from [wu2023unleashing].
We also compare against 3D Diffusion Policy [Ze2024DP3] and ChainedDiffuser [xian2023chaineddiffuser], which we train both on the language annotated training set. We evaluate 3D Diffusion Policy, ChainedDiffuser and 3D Diffuser Actor with the final checkpoints across 3 seeds. We report both the mean and standard deviation of evaluation results. We devise an algorithm to extract keyposes on CALVIN, since prior works do not use keyposes. We define keyposes as those at frames with significant motion change. Both ChainedDiffuser and 3D Diffuser Actor segment the demonstations based on the keyposes. Notably, though keyposes extracted in RLBench have clear structure since they correspond to programmatically labeled 3D waypoints [james2020rlbench, james2022q], keyposes extracted in CALVIN are instead noisy and random, as the benchmark consists of human play trajectories.
For evaluation, prior works [mees2022hulc, black2023zero, li2023vision] predict a maximum temporal horizon of 360 actions to complete each task, while, on average, ground-truth trajectories are only 60 actions long. Our model predicts trajectories and replans after each trajectory has been executed, instead of after each individual actions. On average, it takes 10 trajectory inference steps to complete a task. We thus allow our model to predict a maximum temporal horizon of 60 trajectories for each task.
Train Task completed in a row episodes 1 2 3 4 5 Avg. Len 3D Diffusion Policy [Ze2024DP3] Lang 28.7±0.4 2.7±0.4 0.0±0.0 0.0±0.0 0.0±0.0 0.31±0.04 MCIL [lynch2020language] All 30.4 1.3 0.2 0.0 0.0 0.31 HULC [mees2022hulc] All 41.8 16.5 5.7 1.9 1.1 0.67 RT-1 [brohan2022rt] Lang 53.3 22.2 9.4 3.8 1.3 0.90 ChainedDiffuser [xian2023chaineddiffuser] (60 keyposes) Lang 49.9±0.01 21.1±0.01 8.0±0.01 3.5±0.0 1.5±0.0 0.84±0.02 RoboFlamingo [li2023vision] Lang 82.4 61.9 46.6 33.1 23.5 2.48 SuSIE [black2023zero] All 87.0 69.0 49.0 38.0 26.0 2.69 GR-1 [wu2023unleashing] Lang 85.4 71.2 59.6 49.7 40.1 3.06 3D Diffuser Actor (ours) Lang 93.8±0.01 80.3±0.0 66.2±0.01 53.3±0.02 41.2±0.01 3.35±0.04
Table 4: Zero-shot long-horizon evaluation on CALVIN on 3 random seeds.
We show quantitative results in Table 4. 3D Diffuser Actor outperforms the state-of-the-art. ChainedDiffuser does not work well on this benchmark. The reason is that its deterministic keypose prediction module fails to predict end-effector keyposes accurately due to multimodality present in human demonstrations of CALVIN in comparison to programmatically collected demonstrations of RLBench. Allowing 3D Diffuser Actor to have a longer horizon before terminating increases the performance significantly, which suggests that the model learns to retry under failure.
@@
Table 5: Multi-Task performance on real-world tasks.
4.3 Evaluation in the real world